OpenStudy (anonymous):

Express as a single logarithm. Simplify, if possible. loga a/sqrtx - loga sqrtax (The a's next to log are subscript, a/sqrtx is a fraction, sqrtx means square root of x and the same goes for sqrtax)

OpenStudy (anonymous):

\[\large \log_a(\sqrt{x})-\log_a(\sqrt{ax})\]?

OpenStudy (anonymous):

\[\large \log_a(\frac{a}{\sqrt{x}})-\log_a(\sqrt{ax})\] maybe?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

there are several ways to do this one is to break apart right away, the other is to combine them as \[\log_a(\frac{a}{\sqrt{x}\sqrt{ax}})\]

OpenStudy (anonymous):

or \[\log_a(\frac{a}{x\sqrt{a}})\]

OpenStudy (anonymous):

then since \(\frac{a}{\sqrt{a}}=\sqrt{a}\) you have \[\log_a(\frac{\sqrt{a}}{x})\]

OpenStudy (anonymous):

or \[\log_a(\sqrt{a})-\log_a(x)=\frac{1}{2}-\log_a(x)\]

OpenStudy (anonymous):

thanks satellite, you are helping my daughter so much

OpenStudy (anonymous):

yw